# Rapid Dissipative Ground State Preparation at Chemical Transition States

### v1

@@ III.3 Cost Analysis
Figure 4: Quantum logical resource estimates for a single dissipative time step operator $W(\sqrt{\tau})\approx e^{\tau\mathcal{L}}$ along with the cost of doing QPE across benchmark chemical systems [44]. The resource estimates are calculated from resource estimates provided in [44] for QPE and the block encoding $U_{H}$ using the “DFTHC-BLISS-SA” SOTA electronic structure compression method to minimize TT/Toffoli gate counts. Each point corresponds to a row in Table 2, with Toffoli count on the vertical axis and logical qubits on the horizontal axis. Regions of current and projected hardware capabilities are shaded to illustrate which systems and algorithms may be accessible on near- or medium-term fault-tolerant devices [45].
To assess the practical prospects for efficient dissipative ground-state preparation along reaction paths, we estimate the logical quantum resources required to prepare ground states for strongly correlated chemical systems along such paths, using optimized quantum circuit encodings [44].
Several quantum algorithms can be used to implement time steps that approximate the purely cooling dynamics generated by $\{K_{a}\}_{a\in\mathcal{A}}$. Broadly, these approaches fall into two classes: collision-based schemes, in which interactions with one or more ancilla qubits induce dissipative kicks that drive the system’s energy downward [20, 21]; and quantum operator Fourier transform–based circuits, which realize the desired filter function using coherent ancilla states of the form $\sum_{i}\sqrt{f(t_{i})}\lvert i\rangle$ together with controlled time evolution [46, 22].
Dissipative kicks are considered to be the most viable near-term approach for implementing dissipative dynamics on quantum computers [21]. For this class of methods, the circuit depth required to implement a short-time dissipative evolution is determined by the leakage properties of the chosen filter function.
For the purpose of resource estimation, we select a periodic square-wave filter consisting of $N_{\omega}$ Fourier modes and perform time evolution $\exp(\pm iHt)$ using the QSP/QSVT framework [47, 48, 49, 50, 51]. Then, the cost of implementing a single time step via the $W(\sqrt{\tau})$ operator can be quantified by the total number of calls $N_{U_{H}}$ to the block encoding $U_{H}$ of the electronic Hamiltonian of interest
@@
Reducing the cost of implementing block encodings of electronic structure Hamiltonians $\mathrm{Cost}(U_{H})$ has been a pervasive subject of recent quantum algorithms research. There are now many schemes for compressing and encoding electronic structure Hamiltonians into quantum circuits [30, 44]. At present, the state-of-the-art is achieved by using a combination of several compression protocols resulting in the “DFTHC-BLISS-SA” block encoding [44]. These optimized block encodings have established the lowest quantum resource cost for investigating chemical reactions, enabling quantum solutions to several application areas that carry significant scientific utility.
Fig. 4 summarizes our logical resource estimates for a single dissipative time step $W(\sqrt{\tau})$ along with the cost for quantum phase estimation (QPE) with chemical accuracy $\sigma_{\text{PEA}}=1$ mHa for representative systems for each of these application areas. In the strongly correlated regimes in which these systems live, we report resource estimates for $N_{\omega}=1000$ Fourier modes for the filter function $f(t)$ given that spectral gaps on the order of 1 mHa are common when studying transition states.
We observe that the cost of implementing the dissipative cooling operator $W$ is more than an order of magnitude greater than that of QPE. This motivates the exploration of methods to decrease the total number of $W$ operators that need to be applied at once during any single coherent run of a quantum circuit during the execution of our algorithm. We can employ techniques developed in [52]. The authors demonstrate how machine learning can be used to learn a function $\nu(\mathbf{R};\bm{\gamma})$, where $\bm{\gamma}$ are the neural network parameters and $\mathbf{R}$ are the chemical geometries along the reaction path. This function determines the parameterization of a variational circuit $U(\nu(\mathbf{R};\bm{\gamma}))|\psi_{0}(\mathbf{R})\rangle$, where $|\psi_{0}(\mathbf{R})\rangle$ is some geometry dependent initial state that gets us into the right region of Hilbert space in order to find a low-depth circuit representation $U$ of the ground state at each point along the reaction path. Using this approach to obtain ground states along a reaction path, we could manage with as few as just a single $W$. Our numerical study of the H4 near its square geometry shows that in regions further from the transition state (emulated by the square geometry of H4), only a single $W$ is required to maintain overlap.
@@ IV Discussion
In this work, we develop an efficient quantum algorithm– dissipative evolution–that transports approximate ground states along reaction paths starting from equilibrium chemical geometries (products and reactants) and ending in the classically difficult regime near transition state chemical geometries where we find quantum advantage. Our algorithm incurs a polynomial overhead bounded by $\widetilde{O}(N_{o}^{3})$. If our reaction pathway is sufficiently smooth (smaller values for the Davis-Kahan constant $C_{\text{DK}}$ then we can expect nearly linear polynomial overhead introduced by our algorithm i.e., $\widetilde{O}(C_{\text{DK}}^{2}N_{o}/\epsilon_{E})$ Theorem 3. This runtime improves upon several other quantum algorithms capable of preparing ground states along reaction pathways, including digital adiabatic evolution and phase randomization [2, 31]. Additionally, we validate our algorithm by numerically preparing the ground state of the square H4 molecule whose geometry introduces strong correlation and therefore is a popular benchmark for dissipative ground state preparation algorithms [20, 29]. We conclude with resource estimates for chemical systems that carry significant scientific utility [44]; our resource estimates are largely derived from the analysis of the cooling primitives introduced in [20] with modifications to the filter function and physical regime.
@@
Our results build on recent progress in quantum algorithms for simulating dissipative dynamics, and we provide an end-to-end complexity guarantee under an ETH-motivated rapid-mixing condition for the engineered cooling primitive, implying that even classically hard transition-state instances can be refined efficiently toward the ground state by dissipative evolution. Strongly correlated, classically hard instances of this type are ubiquitous in nature [26, 35, 53, 54]. FeMoco, for example, has long served as a canonical strongly correlated target for quantum chemistry; recent work has demonstrated an equilibrium-geometry classical ansatz with very high overlap for FeMoco, rendering QPE efficient at or near equilibrium [4] and strengthening the case for concentrating quantum resources on out-of-equilibrium regimes where classical tensor-network methods remain most challenged. In our setting, equilibrium-quality classical solutions provide warm starts near reactant or intermediate geometries, while the quantum dissipative step is used to propagate through the transition-state window where classical methods fail, enabling chemical-accuracy barrier heights relevant for screening and design of FeMoco-inspired synthetic analogues [55].
@@
Looking ahead, several complementary directions could further reduce costs and broaden applicability. First, trajectory-style approaches suggest that even faster mixing may be achievable by interleaving dissipative evolution with QPE-based primitives [56]; consistent with this, our logical resource estimates (Fig. 4) indicate that QPE can be substantially cheaper than a single application of $W$, suggesting that QPE interleaving could reduce the required number of $W$ applications. Second, approximate ground-state generation methods such as that of Ceroni et al. [52] may further reduce the total $W$ budget over an entire pathway. Third, it is useful to distinguish basis-set error from algorithmic hardness: O’Gorman et al. [23] prove that approximating the ground-state energy of the electronic-structure Hamiltonian can remain QMA-complete even when restricted to a fixed single-particle basis and fixed particle number, by encoding hardness in the choice of orbitals used to discretize the problem, and they further show NP-completeness of determining the lowest-energy Slater determinant (Hartree–Fock) within such a fixed basis. They also explicitly pose as open directions whether the problem becomes easier when one promises a bounded basis-set error (i.e., that the chosen basis contains a state sufficiently close to the true ground state) and, separately, how complexity changes when the basis is drawn from a parameterized family and must be selected as part of the instance.
@@
From the perspective of reaction-path algorithms, this suggests a principled interface between representation and dynamics: rather than viewing basis selection as a purely classical preprocessing step, one can treat it as a controllable design degree of freedom coupled to the cooling dynamics, with the aim of compressing entanglement growth and shrinking the width of the hard transition-state window before invoking expensive dissipative refinement. More broadly, the nuclear geometry $\mathbf{R}(s)$ should be viewed as an additional design space rather than a fixed linear interpolation: one can adaptively choose the path discretization and even deform the path in $\mathbf{R}$-space to reduce the onset of strong correlation and minimize the total number of costly refinement steps. Finally, since coherent transport, QPE refinement, and dissipative evolution are not mutually exclusive, a unified theory of when coherent transport suffices and when it is optimal to switch to dissipative refinement would be valuable; in this direction, quantum speed-limit results for adiabatic algorithms at zero and finite temperature [57] may provide principled criteria for strategy switching, including variants that transport a Gibbs state and cool only near the end, or that learn finite temperature reaction rates and extrapolate toward the near-room-temperature regime rather than strictly preparing the ground state directly.
@@
Another promising direction for future work is to incorporate into our algorithm the Dissipative Quantum Eigensolver (DQE) framework of Cubitt [58], which formalizes ground-state preparation as an iterated quantum instrument: the algorithm repeatedly applies weak, local generalized measurements (realizing an AGSP as one Kraus branch) and uses an explicit stopping rule that depends only on the observed measurement record, thereby providing an intrinsic operational “success certificate” without requiring coherent phase estimation or deep reversible arithmetic. Notably, Cubitt emphasizes that a large portion of the DQE circuit can be implemented with Clifford operations, with only a single non-Clifford single-qubit rotation applied on a fixed ancilla line in a representative construction—an architectural feature that directly aligns with our emphasis on low-$T$ overhead primitives and mid-circuit measurement as an algorithmic resource. Building on this, Purcell, Rajput, and Cubitt [59] prove that for stabilizer-encoded (geometrically local) Hamiltonians of the type that naturally arise in fermion-to-qubit encodings, a modified DQE procedure under circuit-level depolarizing noise can suppress the additive error in the final ground-space overlap exponentially in the code distance, moving toward fault-tolerant-quality state preparation without the full overhead of universal fault tolerance; in contrast, they show that dissipative quantum computation (as a model for general computation) does not enjoy such an intrinsic robustness advantage over the standard circuit model.
@@ B.2 Lemma 2 & Lemma 3 (Formal): Rapid cooling via Eigenstate Thermalization Hypothesis (ETH)
where $E=(E_{i}+E_{j})/2$, $\bar{A}(E)$ is a smooth microcanonical (diagonal) expectation, $S(E)$ is the microcanonical entropy, $f_{A}(E,\omega)$ is a smooth envelope that varies only on $O(1)$ energy scales, and $R_{jk}$ is an $O(1)$ zero-mean random variable encoding state-to-state fluctuations [60]. Since $e^{S(E_{\star})}\sim D_{\mathrm{shell}}$ (up to subextensive constants set by $\Delta E$ [61]), for $j\neq k$ in the shell and for $\omega$ restricted to a bounded interval one has the scaling
@@
is controlled primarily by how many downhill targets satisfy this inequality. Equivalently, define the success set $S_{j}=\{k:\ E_{k}\leq E_{j}-\Delta E_{\mathrm{layer}}\}$. ETH suggests the approximation [61]
@@
For chemicals in nature, we observe $M=O(N_{o})$ [30]. Therefore, for $\delta_{k}\coloneqq|s_{k}-s_{k-1}|$ we have
@@
We now combine Lemma 4 (geometric Davis-Kahan control of ground-state drift) and Lemmas 5 and 6 to bound the total number of dissipative blocks required to follow a reaction path, and thereby overall gate complexity of this algorithm. Choose a discretization $0=s_{0}<s_{1}<\cdots<s_{N_{H}}=1$ of the reaction coordinate. Starting from an initial warm start $\rho_{0}$ at $s_{0}$, for each $i=0,\dots,N_{H}-1$ apply the dissipative channel at geometry $s_{i}$ for $t_{i}$ steps to obtain $\rho_{i+1}$, which is then used as a warm start for the next geometry $s_{i+1}$. The goal is to achieve a bounded energy error $|\mathrm{tr}(H(1)\rho_{N_{H}})-E_{0}(1)|\leq\epsilon_{E}$ for some $\epsilon_{E}>0$. Choose a discretization $0=s_{0}<s_{1}<\cdots<s_{N_{H}}=1$. Define the resolved geometric density
@@
When warm-start spectral extensivity condition holds uniformly we have that $\overline{L}_{\mu}=O(N_{o})$ (Lemma 6), and $p_{\min}=\Theta(1)$. Simulating dissipative dynamics for evolution time $t$ on quantum computers requires $\widetilde{O}(t)$ logical quantum resources [46]. Therefore, the total algorithmic runtime (gate) complexity is

### v2

@@ III.3 Cost Analysis
To assess the practical prospects for efficient dissipative ground-state preparation along reaction paths, we estimate the logical quantum resources required to prepare ground states for strongly correlated chemical systems along such paths, using optimized quantum circuit encodings [Low_2025].
Several quantum algorithms can be used to implement time steps that approximate the purely cooling dynamics generated by $\{K_{a}\}_{a\in\mathcal{A}}$. Broadly, these approaches fall into two classes: collision-based schemes, in which interactions with one or more ancilla qubits induce dissipative kicks that drive the system’s energy downward [ding2024single, ding2025End]; and quantum operator Fourier transform–based circuits, which realize the desired filter function using coherent ancilla states of the form $\sum_{i}\sqrt{f(t_{i})}\lvert i\rangle$ together with controlled time evolution [chen2023quantum, Chen2025].
Dissipative kicks are considered to be the most viable near-term approach for implementing dissipative dynamics on quantum computers [ding2025End]. For this class of methods, the circuit depth required to implement a short-time dissipative evolution is determined by the leakage properties of the chosen filter function.
For the purpose of resource estimation, we select a periodic square-wave filter consisting of $N_{\omega}$ Fourier modes and perform time evolution $\exp(\pm iHt)$ using the QSP/QSVT framework [gilyen2019quantum, low2017optimal, Low2019hamiltonian, low2016resonant, Haah2019product]. Then, the cost of implementing a single time step via the $W(\sqrt{\tau})$ operator can be quantified by the total number of calls $N_{U_{H}}$ to the block encoding $U_{H}$ of the electronic Hamiltonian of interest
@@
Reducing the cost of implementing block encodings of electronic structure Hamiltonians $\mathrm{Cost}(U_{H})$ has been a pervasive subject of recent quantum algorithms research. There are now many schemes for compressing and encoding electronic structure Hamiltonians into quantum circuits [Motta_2021, Low_2025]. At present, the state-of-the-art is achieved by using a combination of several compression protocols resulting in the “DFTHC-BLISS-SA” block encoding [Low_2025]. These optimized block encodings have established the lowest quantum resource cost for investigating chemical reactions, enabling quantum solutions to several application areas that carry significant scientific utility.
Fig. 6 summarizes our logical resource estimates for a single dissipative time step $W(\sqrt{\tau})$ along with the cost for quantum phase estimation (QPE) with chemical accuracy $\sigma_{\text{PEA}}=1$ mHa for representative systems for each of these application areas. In the strongly correlated regimes in which these systems live, we report resource estimates for $N_{\omega}=1000$ Fourier modes for the filter function $f(t)$ given that spectral gaps on the order of 1 mHa are common when studying transition states. Note that the QPE costs presented are single-shot QPE, and do not account for infidelity of the initial state. There is no way to guarantee (even an inverse polynomial) overlap of the initial ansätz state at the TS configuration, resulting in many repetitions that increase the cost of the QPE, potentially by orders of magnitude that will dwarf the apparent mismatch.
We observe that the cost of implementing the dissipative cooling operator $W$ is more than an order of magnitude greater than that of QPE. This motivates the exploration of methods to decrease the total number of $W$ operators that need to be applied at once during any single coherent run of a quantum circuit during the execution of our algorithm. We can employ techniques developed in [ceroni2023generatingapproximategroundstates]. The authors demonstrate how machine learning can be used to learn a function $\nu(\mathbf{R};\bm{\gamma})$, where $\bm{\gamma}$ are the neural network parameters and $\mathbf{R}$ are the chemical geometries along the reaction path. This function determines the parameterization of a variational circuit $U(\nu(\mathbf{R};\bm{\gamma}))|\psi_{0}(\mathbf{R})\rangle$, where $|\psi_{0}(\mathbf{R})\rangle$ is some geometry dependent initial state that gets us into the right region of Hilbert space in order to find a low-depth circuit representation $U$ of the ground state at each point along the reaction path. Using this approach to obtain ground states along a reaction path, we could manage with as few as just a single $W$. Our numerical study of the H4 near its square geometry shows that in regions further from the transition state (emulated by the square geometry of H4), only a single $W$ is required to maintain overlap.
@@ IV Discussion
In this work, we develop an efficient quantum algorithm– dissipative evolution–that transports approximate ground states along reaction paths starting from equilibrium chemical geometries (products and reactants) and ending in the classically difficult regime near transition state chemical geometries where we find quantum advantage. Our algorithm incurs a polynomial overhead bounded by $\widetilde{O}(N_{o}^{3})$. If our reaction pathway is sufficiently smooth (smaller values for the Davis-Kahan constant $C_{\text{DK}}$ then we can expect nearly linear polynomial overhead introduced by our algorithm i.e., $\widetilde{O}(C_{\text{DK}}^{2}N_{o}/\epsilon_{E})$ Theorem 3. Furthermore, a path can be optimized to maintain smoothness and the gap. This runtime improves upon several other quantum algorithms capable of preparing ground states along reaction pathways, including digital adiabatic evolution and phase randomization [Berry2025Large, Boixo2009Eigenpath]. Additionally, we validate our algorithm by numerically preparing the ground state of the square H4 molecule whose geometry introduces strong correlation and therefore is a popular benchmark for dissipative ground state preparation algorithms [ding2024single, Li_2025]. We conclude with resource estimates for chemical systems that carry significant scientific utility [Low_2025]; our resource estimates are largely derived from the analysis of the cooling primitives introduced in [ding2024single] with modifications to the filter function and physical regime.
@@
Our results build on recent progress in quantum algorithms for simulating dissipative dynamics, and we provide an end-to-end complexity guarantee under an ETH-motivated rapid-mixing condition for the engineered cooling primitive, implying that even classically hard transition-state instances can be refined efficiently toward the ground state by dissipative evolution. Strongly correlated, classically hard instances of this type are ubiquitous in nature [Reiher_2017, Wardzala2024, Rudshteyn2020AFQMC_TMComplexes, otten2024quantumresourcesrequiredbinding]. FeMoco, for example, has long served as a canonical strongly correlated target for quantum chemistry; recent work has demonstrated an equilibrium-geometry classical ansatz with very high overlap for FeMoco, rendering QPE efficient at or near equilibrium [zhai2026classicalsolutionfemoco] and strengthening the case for concentrating quantum resources on out-of-equilibrium regimes where classical tensor-network methods remain most challenged. In our setting, equilibrium-quality classical solutions provide warm starts near reactant or intermediate geometries, while the quantum dissipative step is used to propagate through the transition-state window where classical methods fail, enabling chemical-accuracy barrier heights relevant for screening and design of FeMoco-inspired synthetic analogues [Xu2025-femoco].
@@
Looking ahead, several complementary directions could further reduce costs and broaden applicability. First, trajectory-style approaches suggest that even faster mixing may be achievable by interleaving dissipative evolution with QPE-based primitives [JiangIPAM2026Trajectory]; consistent with this, our logical resource estimates (Fig. 6) indicate that QPE can be substantially cheaper than a single application of $W$, suggesting that QPE interleaving could reduce the required number of $W$ applications. Second, approximate ground-state generation methods such as that of Ceroni et al. [ceroni2023generatingapproximategroundstates] may further reduce the total $W$ budget over an entire pathway. Third, it is useful to distinguish basis-set error from algorithmic hardness: O’Gorman et al. [OGorman2022Intractability] prove that approximating the ground-state energy of the electronic-structure Hamiltonian can remain QMA-complete even when restricted to a fixed single-particle basis and fixed particle number, by encoding hardness in the choice of orbitals used to discretize the problem, and they further show NP-completeness of determining the lowest-energy Slater determinant (Hartree–Fock) within such a fixed basis. They also explicitly pose as open directions whether the problem becomes easier when one promises a bounded basis-set error (i.e., that the chosen basis contains a state sufficiently close to the true ground state) and, separately, how complexity changes when the basis is drawn from a parameterized family and must be selected as part of the instance.
@@
From the perspective of reaction-path algorithms, this suggests a principled interface between representation and dynamics: rather than viewing basis selection as a purely classical preprocessing step, one can treat it as a controllable design degree of freedom coupled to the cooling dynamics, with the aim of compressing entanglement growth and shrinking the width of the hard transition-state window before invoking expensive dissipative refinement. More broadly, the nuclear geometry $\mathbf{R}(s)$ should be viewed as an additional design space rather than a fixed linear interpolation: one can adaptively choose the path discretization and even deform the path in $\mathbf{R}$-space to reduce the onset of strong correlation and minimize the total number of costly refinement steps. Finally, since coherent transport, QPE refinement, and dissipative evolution are not mutually exclusive, a unified theory of when coherent transport suffices and when it is optimal to switch to dissipative refinement would be valuable; in this direction, quantum speed-limit results for adiabatic algorithms at zero and finite temperature [chou2026universalscalingfinitetemperaturequantum] may provide principled criteria for strategy switching, including variants that transport a Gibbs state and cool only near the end, or that learn finite temperature reaction rates and extrapolate toward the near-room-temperature regime rather than strictly preparing the ground state directly.
@@
Another promising direction for future work is to incorporate into our algorithm the Dissipative Quantum Eigensolver (DQE) framework of Cubitt [cubitt2023dissipativegroundstatepreparation], which formalizes ground-state preparation as an iterated quantum instrument: the algorithm repeatedly applies weak, local generalized measurements (realizing an AGSP as one Kraus branch) and uses an explicit stopping rule that depends only on the observed measurement record, thereby providing an intrinsic operational “success certificate” without requiring coherent phase estimation or deep reversible arithmetic. Notably, Cubitt emphasizes that a large portion of the DQE circuit can be implemented with Clifford operations, with only a single non-Clifford single-qubit rotation applied on a fixed ancilla line in a representative construction—an architectural feature that directly aligns with our emphasis on low-$T$ overhead primitives and mid-circuit measurement as an algorithmic resource. Building on this, Purcell, Rajput, and Cubitt [purcell2025faultresiliencedissipativeprocessesquantum] prove that for stabilizer-encoded (geometrically local) Hamiltonians of the type that naturally arise in fermion-to-qubit encodings, a modified DQE procedure under circuit-level depolarizing noise can suppress the additive error in the final ground-space overlap exponentially in the code distance, moving toward fault-tolerant-quality state preparation without the full overhead of universal fault tolerance; in contrast, they show that dissipative quantum computation (as a model for general computation) does not enjoy such an intrinsic robustness advantage over the standard circuit model.
@@ B.2 Lemma 2 & Lemma 3 (Formal): Rapid cooling via Eigenstate Thermalization Hypothesis (ETH)
where $E=(E_{i}+E_{j})/2$, $\bar{A}(E)$ is a smooth microcanonical (diagonal) expectation, $S(E)$ is the microcanonical entropy, $f_{A}(E,\omega)$ is a smooth envelope that varies only on $O(1)$ energy scales, and $R_{jk}$ is an $O(1)$ zero-mean random variable encoding state-to-state fluctuations [Deutsch_2018]. Since $e^{S(E_{\star})}\sim D_{\mathrm{shell}}$ (up to subextensive constants set by $\Delta E$ [DAlessio_2016]), for $j\neq k$ in the shell and for $\omega$ restricted to a bounded interval one has the scaling
@@
is controlled primarily by how many downhill targets satisfy this inequality. Equivalently, define the success set $S_{j}=\{k:\ E_{k}\leq E_{j}-\Delta E_{\mathrm{layer}}\}$. ETH suggests the approximation [DAlessio_2016]
@@
For chemicals in nature, we observe $M=O(N_{o})$ [Motta_2021]. Therefore, for $\delta_{k}\coloneqq|s_{k}-s_{k-1}|$ we have
@@
We now combine Lemma 5 (geometric Davis-Kahan control of ground-state drift) and Lemmas 6 and 7 to bound the total number of dissipative blocks required to follow a reaction path, and thereby overall gate complexity of this algorithm. Choose a discretization $0=s_{0}<s_{1}<\cdots<s_{N_{H}}=1$ of the reaction coordinate. Starting from an initial warm start $\rho_{0}$ at $s_{0}$, for each $i=0,\dots,N_{H}-1$ apply the dissipative channel at geometry $s_{i}$ for $t_{i}$ steps to obtain $\rho_{i+1}$, which is then used as a warm start for the next geometry $s_{i+1}$. The goal is to achieve a bounded energy error $|\mathrm{tr}(H(1)\rho_{N_{H}})-E_{0}(1)|\leq\epsilon_{E}$ for some $\epsilon_{E}>0$. Choose a discretization $0=s_{0}<s_{1}<\cdots<s_{N_{H}}=1$. Define the resolved geometric density
@@
When warm-start spectral extensivity condition holds uniformly we have that $\overline{L}_{\mu}=O(N_{o})$ (Lemma 7), and $p_{\min}=\Theta(1)$. Simulating dissipative dynamics for evolution time $t$ on quantum computers requires $\widetilde{O}(t)$ logical quantum resources [chen2023quantum]. Therefore, the total algorithmic runtime (gate) complexity is